i would really like someone to help me with this problem i have tried to do it but im not having much luck...
NAme the slope of the line and the y intercept
2x+3y+32=0

i tried to subtract 2x from everything but i came out with 1xy+32=0 and that didnt make sense then i tried to change the signs and i got something else funny looking but i really need someone to explain it to me better than my teacher she is a good teacher but i dont understand her....
thanks

Answer by checkley71(8403)

THE EQUATION FOR THE SLOPE INTERCEPT IS Y=mX+b OR IN THIS CASE WE HAVE
2X+3Y+32=0 OR 3Y=-2X-32 OR Y=-2X/3-32/3 THUS THE SLOPE (m)IS -2/3 &NTHE Y INTERCEPT (b) IS -32/3.